# Coherent Information Phase Transition in a Noisy Quantum Circuit

### v2

@@ Abstract
The proxy $\chi$ resembles the “quantum” cross entropy introduced in Ref. [53], with the notable distinction that ancilla qubits are introduced dynamically in this context. Furthermore, while this proxy does not reveal MIPT, it proves to be a robust indicator of the coherent information transition. As demonstrated in Supplemental Material [59], $\chi$, akin to the coherent information, is directly related to the free energy difference between distinct boundary conditions imposed on the bottom boundary. This dependency drives a sharp transition in $\chi$ from 00 to $1$. Physically, this probe reflects the distinguishability of different initial states by leveraging access to both the measurement outcomes and the final state of the ancilla qubits. In the recoverable phase, the initial information is preserved, and we therefore expect $\chi=0$, corresponding to perfect distinguishability. Conversely, in the irrecoverable phase, the initial information becomes erased, causing $\chi$ to approach 1.
To demonstrate the efficacy of this probe, we present numerical results in Fig. 4(b) and Fig. 4(c) for the case of depolarizing noise with $p=0$ and $p=0.1$, respectively. The critical points are determined to be $q_{c}=0.362(2)$ and $q_{c}=0.378(2)$, which are in agreement with the exact phase transition points.
@@
Discussions. The coherent information is related to the single-shot quantum channel capacity by the relation $\mathcal{Q}_{1}=\max_{\rho_{SR}}I_{C}(R>SA)$ [9, 2]. Here we consider $\rho_{SR}$ to be a direct product of Bell pairs between $S$ and $R$. Through analytical mapping, it is evident that this choice maximizes the coherent information within the recoverable phase, minimizing the free energy $\mathcal{F}^{(n,k)}_{SAR}$. However, in the irrecoverable phase, selecting $\rho_{SR}$ as a trivial product state would yield zero coherent information, thereby maximizing the coherent information. This results in a phase transition where the single-shot quantum channel capacity shifts from a finite positive value to zero. Furthermore, the true quantum channel capacity $\mathcal{Q}$, which allows for multiple uses of the channel, becomes additive and equal to $\mathcal{Q}_{1}$ only if the channel is degradable [74, 75, 76, 77, 78, 79, 80]. Although measurement and QE operations are all degradable, general noises often lacks this property, making the calculation of $\mathcal{Q}$ more challenging. Thus, it is intriguing to explore whether a phase transition in $\mathcal{Q}$ can be identified within this framework.
[…]
@@ Acknowledgements.
Acknowledgment. We are deeply indebted to Xiao-Liang Qi for many valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01.

### v3

@@ Abstract
where $\mathbb{E}_{\mathcal{C}}$ represents averaging over different circuit realizations. It resembles the “quantum” cross entropy introduced in Ref. [54], with the notable distinction that ancilla qubits introduced dynamically during the evolution, rather than being present from the beginning. Furthermore, while $\chi$ does not reveal MIPT, it is a robust indicator of the coherent information transition. Intuitively, this probe reflects the distinguishability of different initial states by leveraging access to both the measurement outcomes and the final state of the ancilla qubits. In the recoverable phase, the initial information is preserved, and therefore we expect $\chi=0$, corresponding to perfect distinguishability. In contrast, in the irrecoverable phase, the initial information is erased, causing $\chi$ to approach 1. A rigorous analytical analysis demonstrating the sharp transition of $\chi$ at the phase transition is provided in [60].
Figure 4: Efficient protocol. (a) Schematic workflow of the protocol. An important additional ingredient, compared to the approach in Ref. [54], is the incorporation of projective measurements on the ancilla qubits. (b) Take depolarizing noise as an example. We choose $\rho=(\left|+\right\rangle\left\langle+\right\rangle)^{\otimes L/2}\otimes(\left|0\right\rangle\left\langle 0\right\rangle)^{\otimes L/2}$ and $\sigma=(\left|0\right\rangle\left\langle 0\right\rangle)^{\otimes L}$. Every data point is averaged over $3\times 10^{3}$ realizations. The inset shows the data collapse result. (c) Analogous to (b), with measurement probability $p=0.1$.
A key advantage of $\chi$ is that its measurement allows us to fully utilize each circuit realization and trajectory, circumventing the need for experimental control over noise locations or post-selection on measurement outcomes. The central idea is that the run with initial state $\sigma$ can be classically simulated by selecting it as a stabilizer state. When the unitary gates are Clifford gates, $\widetilde{\sigma_{A}^{m}}$ remains a stabilizer state, leading to the following expression for $\chi$:
$\chi=\mathbb{E}_{\mathcal{C}}\sum_{m}p_{\rho}^{m}\frac{p_{\sigma}^{m}}{2^{-N_{\text{rand}}^{\sigma}}}\text{Tr}\left(\rho_{A}^{m}\prod_{k=1}^{l_{\sigma}}\frac{1+g_{k}}{2}\right).$ (5)
Here, $N_{\text{rand}}^{\sigma}$ is the number of measurements with random outcomes during the classical simulation, $l_{\sigma}$ represents the number of generators for the normalized stabilizer state $\sigma_{A}^{m}$ and $g$ are the corresponding generators. The middle term $p_{\sigma}^{m}/2^{-N_{\text{rand}}^{\sigma}}$ is either 0 or $1$ due to the property of stabilizer formalism, while the final term corresponds to the probability of measuring all the generators with outcomes equal to $1$. Thus, the overall process for measuring $\chi$ is outlined in Fig. 4(a) and is detailed as follows. For each execution of the quantum circuit on a quantum device, we first conduct the classical simulation of the same circuit with initial state $\sigma$, leveraging the knowledge of the measurement outcomes and the circuit realization. We then compute $p_{\sigma}^{m}/2^{-N_{\text{rand}}^{\sigma}}$ via classical simulation. If the result is $1$, we proceed to perform projective measurements on the ancilla qubits in the quantum device according to the stabilizer generators. If all measurement outcomes are $1$, we denote this as a successful event. The quantity $\chi$ is then estimated as the proportion of successful events among all circuit executions. To demonstrate the efficacy of this probe, we present numerical results in Fig. 4(b) and Fig. 4(c) for the case of depolarizing noise with $p=0$ and $p=0.1$, respectively. Here, $\rho$ is chosen as a stabilizer state for simulating the quantum device, but it could in principle be any state, rendering the evolution intractable classically [73]. The critical points are determined to be $q_{c}=0.362(2)$ and $q_{c}=0.378(2)$, which are in agreement with the exact phase transition points.
@@
Discussions—The coherent information is related to the single-shot quantum channel capacity by the relation $\mathcal{Q}_{1}=\max_{\rho_{SR}}I_{C}(R>SA)$ [8, 2]. Here we consider $\rho_{SR}$ to be a direct product of Bell pairs between $S$ and $R$. Through analytical mapping, it is evident that this choice maximizes coherent information within the recoverable phase, minimizing the free energy $\mathcal{F}^{(n,k)}_{SAR}$. However, in the irrecoverable phase, selecting $\rho_{SR}$ as a trivial product state would yield zero coherent information, thereby maximizing the coherent information. This results in a phase transition where the single-shot quantum channel capacity shifts from a finite positive value to zero. Generalizing to the true quantum channel capacity $\mathcal{Q}$, which allows for multiple uses of the channel, general noises are non-degradable, making $\mathcal{Q}$ non-additive and complicating the calculation [74, 75, 76, 77, 78, 79, 80]. Thus, it is intriguing to explore whether a phase transition in $\mathcal{Q}$ can be identified within this framework.
A key aspect of the resource-efficient protocol for detecting the coherent information phase transition is the requirement for knowledge of noise locations. In a practical experimental setting, it has been demonstrated that various types of noise can be effectively converted into erasure errors, which can be located by verifying whether the qubit remains in the computational subspace [81, 82, 83, 84, 85, 86, 87, 88, 89]. Upon detection of an erasure error, replacing the qubit with a maximally mixed state corresponds to a depolarizing noise, while replacing it with $\left|0\right$> corresponds to a resetting noise. It is therefore of interest to investigate whether imperfect detection of erasure errors would compromise the protocol’s success. We leave this to future work.
@@ Acknowledgements.
Acknowledgment. We thank Xiao-Liang Qi and Chong Wang for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01, and is sponsored by the “Shuguang Program” supported by the Shanghai Education Development Foundation and Shanghai Municipal Education Commission.